# HumanEval-XL: A Multilingual Code Generation Benchmark for Cross-lingual Natural Language Generalization

### 26 Feb 2024 v1

@@ Appendix B Comprehensive Experiment Results
English 0.00 0.00 0.00 3.75 11.25 17.50 13.75 0.00 66.25
Russian 0.00 0.00 0.00 7.50 8.75 11.25 10.00 1.25 68.75
Chinese 0.00 0.00 0.00 6.25 10.00 13.75 15.00 0.00 67.50
German 0.00 0.00 0.00 5.00 10.00 6.25 10.00 0.00 72.50
Spanish 0.00 0.00 0.00 6.25 10.00 11.25 12.50 1.25 66.25
French 0.00 0.00 0.00 2.50 8.75 10.00 15.00 0.00 60.00
Italian 0.00 0.00 0.00 7.50 11.25 12.50 13.75 1.25 71.25
Portuguese 0.00 0.00 0.00 6.25 10.00 10.00 13.75 1.25 63.75
Greek 0.00 0.00 0.00 3.75 8.75 7.50 11.25 0.00 61.25
Hungarian 0.00 0.00 0.00 5.00 11.25 8.75 11.25 1.25 65.00
Dutch 0.00 0.00 0.00 5.00 10.00 11.25 15.00 1.25 61.25
Finnish 0.00 0.00 0.00 5.00 7.50 8.75 10.00 1.25 65.00
Indonesian 0.00 0.00 0.00 5.00 10.00 8.75 12.50 0.00 61.25
Turkish 0.00 0.00 0.00 8.75 10.00 7.50 10.00 0.00 51.25
Arabic 0.00 0.00 0.00 2.50 8.75 10.00 8.75 0.00 53.75
Vietnamese 0.00 0.00 0.00 5.00 10.00 10.00 12.50 0.00 47.50
Bulgarian 0.00 0.00 0.00 6.25 7.50 12.50 10.00 0.00 68.75
Persian 0.00 0.00 0.00 2.50 8.75 7.50 11.25 0.00 62.50
Malay 0.00 0.00 0.00 3.75 7.50 6.25 11.25 0.00 63.75
Hebrew 0.00 0.00 0.00 2.50 10.00 8.75 10.00 0.00 60.00
Estonian 0.00 0.00 0.00 3.75 8.75 11.25 12.50 1.25 65.00
Tagalog 0.00 0.00 0.00 5.00 6.25 6.25 12.50 0.00 66.25
Afrikaans 0.00 0.00 0.00 5.00 11.25 8.75 13.75 1.25 68.75
@@
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 6.25 26.25
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 25.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.25 23.75
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 20.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 15.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.25 18.75
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 32.50
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 23.75
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.00 28.75
@@
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.50 41.25
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.50 10.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.25 27.50
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 25.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 27.50
@@
English 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Russian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Chinese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Indonesian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Malay 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hebrew 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Estonian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Tagalog 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Afrikaans 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
@@
English 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Russian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Chinese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Indonesian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Malay 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hebrew 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Estonian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Tagalog 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Afrikaans 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00

### 24 Mar 2024 v2

@@ Appendix B Comprehensive Experiment Results
English 0.00 0.00 0.00 3.75 11.25 17.50 13.75 11.25 66.25
Russian 0.00 0.00 0.00 7.50 8.75 11.25 10.00 17.50 68.75
Chinese 0.00 0.00 0.00 6.25 10.00 13.75 15.00 11.25 67.50
German 0.00 0.00 0.00 5.00 10.00 6.25 10.00 10.00 72.50
Spanish 0.00 0.00 0.00 6.25 10.00 11.25 12.50 11.25 66.25
French 0.00 0.00 0.00 2.50 8.75 10.00 15.00 11.25 60.00
Italian 0.00 0.00 0.00 7.50 11.25 12.50 13.75 12.50 71.25
Portuguese 0.00 0.00 0.00 6.25 10.00 10.00 13.75 12.50 63.75
Greek 0.00 0.00 0.00 3.75 8.75 7.50 11.25 10.00 61.25
Hungarian 0.00 0.00 0.00 5.00 11.25 8.75 11.25 11.25 65.00
Dutch 0.00 0.00 0.00 5.00 10.00 11.25 15.00 11.25 61.25
Finnish 0.00 0.00 0.00 5.00 7.50 8.75 10.00 10.00 65.00
Indonesian 0.00 0.00 0.00 5.00 10.00 8.75 12.50 7.50 61.25
Turkish 0.00 0.00 0.00 8.75 10.00 7.50 10.00 10.00 51.25
Arabic 0.00 0.00 0.00 2.50 8.75 10.00 8.75 13.75 53.75
Vietnamese 0.00 0.00 0.00 5.00 10.00 10.00 12.50 10.00 47.50
Bulgarian 0.00 0.00 0.00 6.25 7.50 12.50 10.00 8.75 68.75
Persian 0.00 0.00 0.00 2.50 8.75 7.50 11.25 8.75 62.50
Malay 0.00 0.00 0.00 3.75 7.50 6.25 11.25 8.75 63.75
Hebrew 0.00 0.00 0.00 2.50 10.00 8.75 10.00 18.75 60.00
Estonian 0.00 0.00 0.00 3.75 8.75 11.25 12.50 15.00 65.00
Tagalog 0.00 0.00 0.00 5.00 6.25 6.25 12.50 6.25 66.25
Afrikaans 0.00 0.00 0.00 5.00 11.25 8.75 13.75 8.75 68.75
@@
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 10.00 26.25
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.00 25.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.75 23.75
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.00 20.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.75 15.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.75 18.75
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.75 32.50
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 10.00 23.75
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.75 28.75
@@
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 6.25 41.25
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 7.50 10.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.50 27.50
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 13.75 25.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 15.00 27.50
@@
English 0.00 0.00 0.00 3.75 6.25 8.75 12.50 3.75 52.50
Russian 0.00 1.25 0.00 2.50 5.00 7.50 12.50 3.75 58.75
Chinese 0.00 1.25 0.00 3.75 5.00 7.50 10.00 1.25 52.50
German 0.00 0.00 0.00 5.00 3.75 5.00 8.75 3.75 55.00
Spanish 0.00 0.00 0.00 3.75 2.50 6.25 11.25 2.50 55.00
French 0.00 0.00 0.00 2.50 5.00 3.75 7.50 0.00 53.75
Italian 0.00 0.00 0.00 2.50 3.75 5.00 12.50 1.25 56.25
Portuguese 0.00 0.00 0.00 3.75 3.75 7.50 8.75 1.25 53.75
Greek 0.00 1.25 0.00 3.75 5.00 6.25 8.75 2.50 51.25
Hungarian 0.00 0.00 0.00 3.75 6.25 6.25 7.50 2.50 57.50
Dutch 0.00 0.00 0.00 3.75 5.00 6.25 10.00 6.25 56.25
Finnish 0.00 0.00 0.00 3.75 6.25 6.25 7.50 0.00 55.00
Indonesian 0.00 0.00 0.00 3.75 5.00 7.50 11.25 0.00 58.75
Turkish 0.00 0.00 0.00 2.50 5.00 5.00 7.50 1.25 57.50
Arabic 0.00 1.25 0.00 2.50 3.75 5.00 6.25 2.50 55.00
Vietnamese 0.00 0.00 0.00 2.50 3.75 6.25 10.00 2.50 56.25
Bulgarian 0.00 1.25 0.00 2.50 2.50 6.25 11.25 1.25 57.50
Persian 0.00 1.25 0.00 2.50 5.00 6.25 8.75 1.25 58.75
Malay 0.00 0.00 0.00 3.75 6.25 6.25 10.00 1.25 55.00
Hebrew 0.00 1.25 0.00 3.75 3.75 5.00 10.00 5.00 55.00
Estonian 0.00 0.00 0.00 2.50 5.00 6.25 6.25 2.50 56.25
Tagalog 0.00 0.00 0.00 3.75 6.25 5.00 6.25 2.50 56.25
Afrikaans 0.00 0.00 0.00 3.75 3.75 5.00 6.25 5.00 57.50
@@
English 0.00 1.25 1.25 7.50 16.25 21.25 22.50 11.25 75.00
Russian 0.00 2.50 0.00 8.75 11.25 18.75 23.75 8.75 70.00
Chinese 0.00 1.25 0.00 6.25 13.75 18.75 23.75 3.75 68.75
German 0.00 1.25 0.00 10.00 13.75 20.00 20.00 3.75 70.00
Spanish 0.00 1.25 0.00 10.00 11.25 23.75 21.25 5.00 73.75
French 0.00 0.00 1.25 8.75 13.75 20.00 21.25 1.25 71.25
Italian 0.00 1.25 0.00 8.75 13.75 16.25 20.00 5.00 72.50
Portuguese 0.00 0.00 0.00 7.50 13.75 20.00 21.25 6.25 71.25
Greek 0.00 3.75 0.00 10.00 12.50 15.00 20.00 5.00 65.00
Hungarian 0.00 0.00 0.00 10.00 12.50 13.75 18.75 5.00 76.25
Dutch 0.00 0.00 0.00 7.50 16.25 20.00 20.00 6.25 70.00
Finnish 0.00 0.00 0.00 10.00 15.00 16.25 20.00 6.25 73.75
Indonesian 0.00 0.00 0.00 6.25 15.00 18.75 20.00 2.50 71.25
Turkish 0.00 2.50 0.00 11.25 12.50 17.50 18.75 2.50 68.75
Arabic 0.00 3.75 0.00 10.00 12.50 17.50 25.00 3.75 66.25
Vietnamese 0.00 1.25 0.00 11.25 15.00 15.00 22.50 6.25 70.00
Bulgarian 0.00 2.50 0.00 10.00 11.25 17.50 21.25 7.50 70.00
Persian 0.00 5.00 0.00 8.75 15.00 22.50 18.75 5.00 72.50
Malay 0.00 1.25 0.00 8.75 12.50 20.00 18.75 6.25 71.25
Hebrew 0.00 2.50 0.00 8.75 12.50 17.50 20.00 5.00 71.25
Estonian 0.00 0.00 0.00 11.25 12.50 16.25 22.50 7.50 72.50
Tagalog 0.00 0.00 0.00 8.75 12.50 17.50 18.75 6.25 72.50
Afrikaans 0.00 0.00 0.00 11.25 13.75 17.50 21.25 1.25 70.00